OpenStudy (anonymous):

A square pyramid has a volume of 20 cubic units. What would be the change to the volume if the dimensions of the square base were doubled? The volume would be twice the original. The volume would be three times the original. The volume would be four times the original. The volume would be eight times the original

OpenStudy (anonymous):

The volume would be twice the original since the base area gets doubled. \[\frac{ 1 }{ 3} \times (Base Area) \times Height\] Now, twice the Base Area would be : \[\frac{ 1 }{ 3} \times (2)(Base Area) \times Height\] So, the entire thing gets doubled on the other side as well. Which means that 20 cubic units becomes 40 cubic units.

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

You're welcome. ^_^

OpenStudy (anonymous):

incorrect.